OpenStudy (bbb911):

This is Geometry: Find the area of an equilateral triangle with an apothem 3.5 meters long and each side 12 meters long.

OpenStudy (anonymous):

heigh = 2*3.5=7

OpenStudy (anonymous):

A= (12*7)/2 A=42m^2

OpenStudy (bbb911):

thanks again for your help :)